### v1

@@ Abstract
Over the past few decades, statistical methods for causal inference have made impressive strides, enabling progress across a range of scientific fields. However, these methodological advances have often been confined to individual studies, limiting their ability to draw more generalizable conclusions. Achieving a thorough understanding of cause and effect typically relies on the integration, reconciliation, and synthesis from diverse study designs and multiple data sources. Furthermore, it is crucial to direct this synthesis effort toward understanding the effect of treatments for specific patient populations. To address these challenges, we present a weighting framework for evidence synthesis that handles both individual- and aggregate-level data, encompassing and extending conventional regression-based meta-analysis methods. We use this approach to tailor meta-analyses, targeting the covariate profiles of patients in a target population in a sample-bounded manner, thereby enhancing their personalization and robustness. We propose a technique to detect studies that meaningfully deviate from the target population, suggesting when it might be prudent to exclude them from the analysis. We establish multiple consistency conditions and demonstrate asymptotic normality for the proposed estimator. We illustrate this approach using simulated and real data.
@@ 1 Introduction
In recent years, the field of causal inference has seen remarkable progress, yet much of this methodological innovation has been confined to isolated studies, each conducted under varying conditions. Differences in study design, enrolled populations, and outcome measurements can obscure the pursued effects, yielding conflicting results. Consequently, meta-analytic techniques have emerged as a key tool to integrate evidence from multiple data sources, often producing more accurate estimates than those of individual studies (Colnet et al. 2024; Rosenbaum 2021; Schmid et al. 2020). The increasing sophistication and adoption of these techniques are evident in the growing literature, underscoring the central role that meta-analysis occupies in assembling credible and generalizable evidence (Figure 1).
@@
In meta-analysis, there are two primary frameworks, defined by the level of granularity of the available data: individual-level data and aggregate-level estimates (Simmonds et al. 2005). Analyses using individual-level data are typically conducted from multiple, separately collected datasets and generally offer greater modeling flexibility and improved statistical efficiency. However, such analyses remain relatively uncommon due to practical and organizational barriers in data acquisition. In contrast, meta-analyses based on aggregate-level data are much more common, but they are inherently constrained by the absence of finely-grained data, which limits the flexibility of statistical adjustments and the scope of effect estimates. In what follows, for clarity and symmetry, we refer to the aggregate-level data framework as AD and the individual-level data framework as ID (although the latter is commonly termed IPD, for individual participant data, in the literature).
@@
ID meta-analyses are typically conducted using either one-stage or two-stage regression methods. The one-stage approach simultaneously analyzes all raw individual-level data using mixed-effects multilevel regression models, while the two-stage approach first derives summary estimates for each study and then combines these in a standard meta-analysis. AD meta-analyses often rely on study-level summary statistics and are commonly implemented through meta-regression, where aggregated covariates are incorporated into the analysis to explore heterogeneity or potential effect modifiers (Tipton et al. 2019).
@@
Despite their widespread use, traditional meta-analytical approaches have some important limitations. First, they do not explicitly reference a well-defined target population and often lack a clear causal interpretation beyond the study samples (Berenfeld et al. 2025; Dahabreh et al. 2020; Vo et al. 2019). The study populations may not reflect the broader target population where treatment decisions are ultimately made, resulting in estimates that do not necessarily apply to the intended group of individuals (Degtiar and Rose 2023). To address this issue, Dahabreh et al. (2020) proposed a framework for causally interpretable meta-analysis to estimate a treatment effect in a specific target population by modeling both treatment assignment and study selection. Building on this work, a growing body of methods has been developed for causally interpretable meta-analysis (Clark et al. 2023; Rott et al. 2024; Steingrimsson et al. 2024). While some recent approaches can combine individual- and aggregate-level data (Rott et al. 2025; Vo et al. 2025), to the best of our knowledge, no existing causally interpretable methods can accommodate the setting where only aggregate-level data are available. Furthermore, most approaches require individual-level data from the target population, which are often unavailable (Dahabreh et al. 2023; Hong et al. 2025; Rott et al. 2025). Finally, it is crucial not only to define a population clearly but also to deliberately select it so that medical recommendations can be tailored to the specific characteristics of patients, enabling personalized healthcare.
@@
Second, effectively and transparently targeting effect estimates to specific patient populations requires a clear understanding of how individual studies or patient measurements contribute to the overall estimate. In particular, data sources that closely align with the target population should be prioritized, highlighting the importance of using estimators that give more weight to relevant information sources and avoid extrapolation beyond the support of the actual measurements. Traditional methods may inadvertently give undue weight to less similar studies, affecting the accuracy and usefulness of the results. Relatedly, issues such as the sign-reversal phenomenon in regression (Small et al. 2017; which can lead to counterintuitive negative weights, where increasing an individual’s outcome under a treatment can decrease the average outcome for that treatment group) underscore the need for transparent diagnostics. Overall, it is desirable for meta-analyses to incorporate diagnostics that quantify the contribution and influence of each study on the final estimate, detect issues such as extrapolation and sign-reversal, and identify which study donors have the greatest weight on the analysis.
@@
To address these challenges, we introduce a weighting framework for evidence synthesis that accommodates both individual- and aggregate-level data, and that both encompasses and extends conventional regression-based meta-analysis methods. This approach allows meta-analyses to be directed to the covariate profiles of target patient populations in a manner that effect estimates are bounded by the observed data, thereby improving both personalization and robustness. We also propose a diagnostic technique for identifying measurements that substantially deviate from the defined target profile, providing guidance on when their exclusion may be warranted. This result is of independent interest in regression analysis, as it demonstrates that, in large samples, regression tends to assign positive weights to observations (or studies) that fall within the support of the target population. In meta-analysis, this diagnostic provides a practical means to detect and trim deviant studies directly from the original data, for example, without the need to estimate propensity scores. Finally, we present a general weighting estimator for meta-analysis, establish multiple consistency conditions, and demonstrate the asymptotic normality of the proposed estimator.
Our work is organized as follows. Section 2 outlines the problem setup, defines a generic target population, and discusses identification with both individual- and aggregate-level data. Section 3 introduces the personalized and sample bounded meta-analysis weighting method, and shows that conventional one-stage and two-stage ID meta analyses are special cases within this framework. Section 4 derives formal properties of the method, including the selection mechanism induced by sample boundedness, multiple consistency conditions, and asymptotic normality. Section 5 details diagnostic procedures and provides theoretical guarantees that the method can identify relevant study donors under correctly specified models. Section 6 evaluates empirical performance through simulation studies, emphasizing robustness when the target and study populations do not fully overlap. Finally, Section 7 presents two case studies, one using ID and the other using AD, to illustrate the application of the proposed method.
@@ 2 Target estimand and identification assumptions
In settings where only AD are available, individual-level covariates, treatments, and outcomes are not observed. Instead, for each study $i$, we observe the estimated causal effect $\widehat{\tau}_{i}$, its variance estimate $\widehat{\sigma}_{i}^{2}$, and the average covariate vector $\overline{X}_{i}$.
Suppose the $n$ individuals constitute a simple random sample from the study population $\mathcal{P}$. Let $\mathcal{T}$ be the target population. Denote $\mathbb{P}$ and $\mathbb{T}$ as the probability measures that characterize $\mathcal{P}$ and $\mathcal{T}$, with respective densities $p(\cdot)$ and $t(\cdot)$. Define the probability measures for the treated and control subpopulations as $\mathbb{P}_{1}=\mathbb{P}_{\cdot\mid Z=1}$ and $\mathbb{P}_{0}=\mathbb{P}_{\cdot\mid Z=0}$, with corresponding densities $p_{1}(\cdot)$ and $p_{0}(\cdot)$. Similarly, put $\mathbb{P}_{(i)}=\mathbb{P}_{\cdot\mid G=i}$ for the measure of the subpopulation of study $i$, with density $p_{(i)}(\cdot)$, for $i=1,\ldots,m$. Our estimand of interest is the average treatment effect for the target population $\mathcal{T}$:
@@
Further, write $\mathcal{Q}=\mathcal{T}\cup\mathcal{P}$ for the combined target and study population, with probability measure $\mathbb{Q}$.
@@ 2.1 Identification with ID
The target average treatment effect $\tau$ is identifiable in both randomized experiments and observational studies, provided the following conditions are satisfied.
@@
Under Assumption 1, the weighting factor is non-negative and satisfies the balancing condition in (2), which guarantees that the distribution of any function of the pretreatment covariates is balanced towards their expectation in the target population. Specifically, setting $f(x)=1$ in (2) demonstrates that the expectation of the weighting factor is one.
@@ 2.2 Identification with AD
When only aggregate-level data are available, identifying the target causal effect becomes more challenging. In this section, we formalize identification in this setting. A common approach in this context is to synthesize evidence by forming a linear combination of study-level estimates. However, the absence of individual-level covariates forces all individuals within a study to share the same weight, limiting the granularity of the adjustments. This constraint hinders our ability to construct weights that accurately align the source studies with the target population.
@@
We now introduce a set of additional assumptions specific to the AD setting. Assumption 2 (a) states that potential outcomes are conditionally exchangeable across studies given the covariates. Part (b) requires the study-level estimators to be unbiased. Part (c) further stipulates the existence of a set of weights that align the study-level covariate distributions with that of the target population. This last condition is particularly stringent and may be difficult to satisfy in practice. In the AD setting, each study $i$ is assigned a single weight $\tilde{w}_{(i)}$ that applies uniformly to all individuals within that study, and this weight cannot adapt to individual-level covariates. As a result, a feasible solution for the weights $\tilde{w}_{(i)}$ may not exist, especially when the number of studies $m$ is small.
@@
Despite the strictness of Assumption 2(c), Proposition 1 establishes that, given conditions (a) and (b), condition (c) is both necessary and sufficient for identification in the case of unrestricted outcome models.
@@ 3 Methods
where $\mathcal{D}(w)$ measures the dispersion of the weights, $\mathcal{B}$ encodes covariate balance constraints, and $\mathcal{A}$ imposes additional adjustment constraints such as normalization and sample-boundedness (Wang and Zubizarreta 2020). Specific estimators arise from different specifications of $\mathcal{D}$, $\mathcal{B}$, and $\mathcal{A}$. We describe these choices in detail in the following sections. A complete summary of the notation is provided in Table S.1 in the Supplementary Materials.
@@ 3.1 One-stage method for ID
where $\widehat{w}^{*+}_{z,ij}$ for $z\in\{0,1\}$ are obtained by solving the following optimization problem:
@@
This balancing constraint operates at two distinct levels. First, across-study balancing ensures that the basis functions in set $A$ are aligned with the target covariate profile across the entire dataset. Second, within-study balancing enforces alignment of the basis functions in set $W$ within each study, addressing potential between-study heterogeneity. Together, these two levels of balancing provide flexibility, allowing the method to effectively address both global and local covariate imbalances. The sets $A$ and $W$ partition the full set of basis functions $\{1,\ldots,K\}$. Personalization is made possible by specifying a target covariate profile as a vector $B^{*}=(B_{1}^{*},\ldots,B_{K}^{*})^{\text{T}}\in\mathbb{R}^{K}$. The adjustment constraint is defined as ${\mathcal{A}}^{+}:=\{\sum_{i,j:Z_{ij}=z}w_{z,ij}=1;\,w_{z,ij}\geq 0\}$, which enforces both normalization and sample-boundedness. Specifically, normalization guarantees that the weights sum to one, while non-negativity prevents extrapolation, ensuring the estimator remains sample-bounded.
@@
The proposed optimization framework (7) offers a means of tailoring inferences for a specific population $\mathcal{T}$ by incorporating the target covariate profile vector $B^{*}$. This profile may correspond to a particular population of interest, thereby facilitating the generalization and transportation of causal inferences. Importantly, balancing toward this profile does not require access to individual-level data in the target population, which may be unavailable for confidentiality reasons. By using summary-level information, such as the sample means of covariates, the proposed framework allows for causally-interpretable inference while maintaining data privacy, making it both flexible and practical for real-world applications where individual-level data may be limited.
@@ 3.2 Two-stage method for AD and mixed data
Our method extends to the mixed data setting, where some or even all studies provide only aggregate-level data. Algorithm 1 outlines a two-stage procedure for this setting.
In the first stage, we perform within-study estimation for each study with available ID using the proposed personalized, sample-bounded approach. Specifically, for each study, we solve the same optimization problem (7) as in the one-stage method, but restrict the analysis to data from that study alone. We then compute the study-specific causal effect estimator $\widehat{\tau}_{i}$ and its corresponding variance estimator $\widehat{\sigma}^{2}_{i}$ using the resulting weights. For variance estimation, we consider two approaches: a heuristic and a plug-in method, with further details provided in Section A.1. The resulting estimates, $\widehat{\tau}_{i}$ and $\widehat{\sigma}^{2}_{i}$, serve as analogues to the aggregate-level statistics available from AD studies. This stage is flexible, allowing each study to include a different set of covariates as permitted by data availability.
@@
Compute the final variance estimator $\widehat{V}^{\mathrm{heur}}_{\mathrm{AD}}$ from (S.3) based on weights

### v2

@@ Abstract
Over the past few decades, statistical methods for causal inference have made impressive strides, enabling progress across a range of scientific fields. However, much of this methodological development has been confined to individual studies, limiting its ability to draw more generalizable conclusions. Achieving a thorough understanding of cause and effect typically relies on the integration, reconciliation, and synthesis from diverse study designs and multiple data sources. Furthermore, it is crucial to direct this synthesis effort toward understanding the effect of treatments for specific patient populations. To address these challenges, we present a weighting framework for evidence synthesis that handles both individual- and aggregate-level data, encompassing and extending conventional regression-based meta-analysis methods. We use this approach to tailor meta-analyses, targeting the covariate profiles of patients in a target population in a sample-bounded manner, thereby enhancing their personalization and robustness. We propose a technique to detect studies that meaningfully deviate from the target population, suggesting when it might be prudent to exclude them from the analysis. We establish multiple consistency conditions and demonstrate asymptotic normality for the proposed estimator. We demonstrate the effectiveness of the method through a simulation study and two real-world case studies.
@@ 1 Introduction
In recent years, the field of causal inference has seen remarkable progress, yet much of this innovation has been confined to isolated studies, each conducted under varying conditions. Differences in study design, enrolled populations, and outcome measurements can obscure treatment effects, yielding conflicting results. Consequently, study integration and meta-analytic techniques have emerged as key tools to synthesize evidence from multiple data sources, often yielding more accurate estimates than those of individual studies (Colnet et al. 2024; Rosenbaum 2021; Schmid et al. 2020). The increasing sophistication and adoption of these techniques are evident in the growing literature, underscoring the central role that meta-analysis plays in assembling credible and generalizable evidence (Figure 1).
@@
In meta-analysis, there are two primary frameworks, defined by the level of granularity of the available data: individual-level and aggregate-level data (Simmonds et al. 2005). Analyses using individual-level data are conducted with multiple, separately collected datasets and generally offer greater modeling flexibility and improved statistical efficiency. However, such analyses remain relatively uncommon due to practical and organizational barriers in data sharing and acquisition. In contrast, meta-analyses based on aggregate-level data are much more common, but they are inherently constrained by the lack of granular information, which limits the flexibility of statistical adjustments and the scope of effect estimates. In what follows, we use the term ID (individual-level data) in place of the more common IPD (individual participant data) terminology to emphasize that our framework is not restricted to patient settings, and to maintain symmetry with the term AD (aggregate-level data).
@@
ID meta-analyses are commonly conducted using either one-stage or two-stage regression methods. The one-stage approach simultaneously analyzes all raw individual-level data using mixed-effects multilevel regression models, while the two-stage approach first derives summary estimates for each study and then combines these in a standard meta-analysis. AD meta-analyses often rely on study-level summary statistics and are commonly implemented through meta-regression, where aggregated covariates are incorporated into the analysis to explore heterogeneity or potential effect modifiers (Tipton et al. 2019).
@@
Despite their widespread use, traditional meta-analytical approaches face several important limitations. First, they often fail to specify a well-defined target population and thus lack a clear causal interpretation beyond the study samples (Berenfeld et al. 2025; Dahabreh et al. 2020; Vo et al. 2019). As a result, study populations may not represent the broader target population where treatment decisions are ultimately made, resulting in estimates that do not necessarily apply to the intended group of individuals (Degtiar and Rose 2023). To address this issue, Dahabreh et al. (2020) proposed a framework for causally interpretable meta-analysis to estimate a treatment effect in a specific target population by modeling both treatment assignment and study selection. Building on this work, a growing body of methods has been developed for causally interpretable meta-analysis (Clark et al. 2023; Rott et al. 2024; Steingrimsson et al. 2024). While some recent approaches can combine individual- and aggregate-level data (Rott et al. 2025; Vo et al. 2025), to our knowledge, no existing causally interpretable methods can accommodate the setting where only aggregate-level data are available. Furthermore, most approaches require individual-level data from the target population, which are often unavailable (Dahabreh et al. 2023; Hong et al. 2025; Rott et al. 2025). Finally, it is crucial not only to define a population clearly, but also to deliberately select it so that medical recommendations can be tailored to the specific characteristics of patients, enabling personalized healthcare.
@@
Second, effectively and transparently targeting effect estimates to specific patient populations requires a clear understanding of how individual studies or patient measurements contribute to the overall estimate. This requires prioritizing data sources that closely align with the target population and using estimators that give more weight to relevant information sources and avoid extrapolation beyond the support of the actual measurements. Traditional methods may inadvertently give undue weight to less similar studies, affecting the accuracy and usefulness of the results. Moreover, known issues such as the sign-reversal phenomenon in regression (Small et al. 2017; which can lead to counterintuitive negative weights, where increasing an individual’s outcome under a treatment can decrease the average outcome for that treatment group) underscore the need for transparent diagnostics. Ideally, a meta-analytic framework should incorporate diagnostics that quantify the contribution and influence of each study on the final estimate, detect issues such as extrapolation and sign-reversal, and identify which “donor” studies have the greatest weight on the analysis.
@@
To address these challenges, we introduce a weighting framework for evidence synthesis that accommodates both individual- and aggregate-level data, and that both encompasses and extends conventional regression-based meta-analysis methods. Our approach enables targeting meta-analyses to the covariate profiles of specific patient populations while ensuring that effect estimates are bounded by the observed data, thereby improving both personalization and robustness. To enhance transparency, we also propose a diagnostic technique for detecting measurements that substantially deviate from the defined target profile, providing guidance on when their exclusion may be warranted. A key theoretical result of our work, which is of independent interest in regression analysis, demonstrates that, in large samples, our method tends to assign positive weights to observations (or studies) that fall within the support of the target population, while systematically excluding those outside it by assigning them zero weight (Theorem 4). In meta-analysis, this diagnostic provides a practical means to detect and trim deviant studies directly from the original data, for example, without the need to estimate propensity scores. Finally, we present a general weighting estimator for meta-analysis, establish its consistency under multiple conditions, and demonstrate its asymptotic normality.
Our work is organized as follows. Section 2 outlines the problem setup, defines a generic target population, and discusses identification with both individual- and aggregate-level data. Section 3 introduces the personalized and sample bounded meta-analysis weighting method, and shows that conventional one-stage and two-stage ID meta analyses are special cases within this framework. Section 4 derives formal properties of the method, including the selection mechanism induced by sample boundedness, multiple consistency conditions, and asymptotic normality. Section 5 details diagnostic procedures and provides theoretical guarantees that the method can identify relevant study donors under correctly specified models. Section 6 evaluates the method’s empirical performance through simulation studies, emphasizing robustness when the target and study populations do not fully overlap. Finally, Section 7 presents two case studies, one using ID and the other using AD, to illustrate the application of the proposed method.
@@ 2 Target estimand and identification assumptions
In settings where only AD are available, individual-level covariates, treatments, and outcomes are not directly observed. Instead, for each study $i$, we observe the estimated causal effect $\widehat{\tau}_{i}$, its variance estimate $\widehat{\sigma}_{i}^{2}$, and the average covariate vector $\overline{X}_{i}$.
Suppose the $n$ individuals constitute a simple random sample from the study population $\mathcal{P}$. Let $\mathcal{T}$ be the target population of interest. Denote $\mathbb{P}$ and $\mathbb{T}$ as the probability measures that characterize $\mathcal{P}$ and $\mathcal{T}$, with respective densities $p(\cdot)$ and $t(\cdot)$. Define the probability measures for the treated and control subpopulations as $\mathbb{P}_{1}=\mathbb{P}_{\cdot\mid Z=1}$ and $\mathbb{P}_{0}=\mathbb{P}_{\cdot\mid Z=0}$, with corresponding densities $p_{1}(\cdot)$ and $p_{0}(\cdot)$. Similarly, put $\mathbb{P}_{(i)}=\mathbb{P}_{\cdot\mid G=i}$ for the measure of the subpopulation of study $i$, with density $p_{(i)}(\cdot)$, for $i=1,\ldots,m$. Our estimand of interest is the average treatment effect for the target population $\mathcal{T}$:
@@
Further, write $\mathcal{Q}=\mathcal{T}\cup\mathcal{P}$ for the combined target and study population, with corresponding probability measure $\mathbb{Q}$.
@@ 2.1 Identification with ID
We can identify the target average treatment effect, $\tau$, from individual-level data (ID) in both randomized experiments and observational studies, provided the following conditions are satisfied.
@@
Under Assumption 1, the weighting factor is non-negative and satisfies the balancing condition in (2), which guarantees that the distribution of any function of the pretreatment covariates is balanced towards its expectation in the target population. Specifically, setting $f(x)=1$ in (2) demonstrates that the expectation of the weighting factor is one.
@@ 2.2 Identification with AD
When only aggregate-level data (AD) are available, identification of the target causal effect becomes more challenging. Here, we formalize the conditions required for identification in this setting. A common approach in this context is to synthesize evidence by forming a linear combination of study-level estimates. However, the absence of individual-level covariates forces all individuals within a study to share the same weight, limiting the granularity of the adjustments. This constraint hinders our ability to construct weights that accurately align the source studies with the target population.
@@
Identification with AD requires several additional assumptions. Assumption 2 (a) states that potential outcomes are conditionally exchangeable across studies given the covariates. Part (b) requires the study-level estimators to be unbiased. Part (c) further stipulates the existence of a set of weights that align the study-level covariate distributions with that of the target population. This last condition is particularly stringent because, in the AD setting, each study $i$ is assigned a single weight $\tilde{w}_{(i)}$ that applies uniformly to all individuals within that study and cannot adapt to individual-level covariates. As a result, a feasible solution for the weights $\tilde{w}_{(i)}$ may not exist, especially when the number of studies $m$ is small.
@@
Proposition 1 shows that, despite its strictness, Assumption 2(c) is not only sufficient but also necessary for identification when the outcome model is unrestricted, provided that parts (a) and (b) of the assumption also hold.
@@ 3 Methods
where $\mathcal{D}(w)$ measures the dispersion of the weights, $\mathcal{B}$ encodes covariate balance constraints, and $\mathcal{A}$ imposes additional adjustment constraints such as normalization and sample-boundedness (Wang and Zubizarreta 2020). Crucially, this formulation allows the weighting factor in Equation 1 to be estimated in a single step, avoiding the multiplicative structure of common multi-stage weighting procedures and yielding more stable estimates (Chattopadhyay et al. 2024). Specific estimators arise from different specifications of $\mathcal{D}$, $\mathcal{B}$, and $\mathcal{A}$. We describe these choices in detail in the following sections. A complete summary of the notation is provided in Table S.1 in the Supplementary Materials.
@@ 3.1 One-stage method for ID
where the weights $\widehat{w}^{*+}_{z,ij}$ for $z\in\{0,1\}$ are obtained by solving the following optimization problem:
@@
This balancing constraint operates at two distinct levels. First, across-study balancing ensures that the basis functions in set $A$ are aligned with the target covariate profile across the entire dataset. Second, within-study balancing enforces alignment of the basis functions in set $W$ within each study, addressing potential between-study heterogeneity. Together, these two levels of balancing provide flexibility, allowing the method to effectively address both global and local covariate imbalances. The sets $A$ and $W$ partition the full set of basis functions $\{1,\ldots,K\}$. Personalization is made possible by specifying a target covariate profile as a vector $B^{*}=(B_{1}^{*},\ldots,B_{K}^{*})^{\text{T}}\in\mathbb{R}^{K}$. The adjustment constraint is defined as ${\mathcal{A}}^{+}:=\{\sum_{i,j:Z_{ij}=z}w_{z,ij}=1;\,w_{z,ij}\geq 0\}$, which enforces both normalization and sample-boundedness. Specifically, normalization guarantees that the weights sum to one, while non-negativity prevents extrapolation, ensuring the estimator remains sample-bounded. In sections 4 and 5, we show that this sample-boundedness constraint translates into a selection mechanism, which helps to detect the data points (e.g., study donors) that are most relevant to the target population.
@@
The proposed optimization framework (7) provides a flexible way to tailor inferences for a specific population $\mathcal{T}$ by incorporating the target covariate profile vector $B^{*}$. This profile may correspond to a particular population of interest, thereby facilitating the generalization and transportation of causal inferences. Crucially, balancing toward this profile does not require access to individual-level data in the target population, which may be unavailable for confidentiality reasons. By using summary-level information, such as the sample means of covariates, the proposed framework allows for causally-interpretable inference while maintaining data privacy, making it both flexible and practical for real-world applications where individual-level data may be limited.
@@ 3.2 Two-stage method for AD and mixed data
Our method also extends to the mixed data setting, where some or even all studies provide only aggregate-level data (AD). Algorithm 1 outlines a two-stage procedure for this setting.
In the first stage, we perform within-study estimation for each study with available ID using the proposed personalized, sample-bounded approach. Specifically, for each study, we solve the same optimization problem (7) as in the one-stage method, but restrict the analysis to data from that study alone. We then compute the study-specific causal effect estimator $\widehat{\tau}_{i}$ and its corresponding variance estimator $\widehat{\sigma}^{2}_{i}$ using the resulting weights. For variance estimation, we consider two approaches: a heuristic and a plug-in method, with further details provided in Section A.1. The resulting estimates, $\widehat{\tau}_{i}$ and $\widehat{\sigma}^{2}_{i}$, serve as analogues to the aggregate-level statistics available from AD studies. A key feature of this stage is its flexibility, as each within-study analysis can use a different set of covariates based on data availability.
@@
Compute the final variance estimator $\widehat{V}^{\mathrm{heur}}_{\mathrm{AD}}$ from (S.3)